Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financial difficulty for some time. The company’s contribution format income statement for the most recent month is given below: Sales (12,800 units × $30 per unit) $ 384,000 Variable expenses 230,400 Contribution margin 153,600 Fixed expenses 171,600 Net operating loss $ (18,000 )

Required: 1. Compute the company’s CM ratio and its break-even point in unit sales and dollar sales.

2. The president believes that a $6,900 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $83,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income?

3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $38,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)?

4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by $0.70 per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,400?

5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $51,000 each month. a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales. b. Assume that the company expects to sell 20,300 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.) c. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,300)?
Answer #1
contribution margin per unit= 153600/12800
12
1) CM ratio = contribution/sales
153600/384,000
40.00%
BEP(units) = total fixed cost/contribution margin per unit
171,600/12
14300
BEP(dollars) = 14300*30
429000
CM ratio 40%
Break even point in units 14300
Break even point in dollars 429000
2) increase in contribution (83000*40%) 33200
less : increase in advertising budget 6,900
increase in net income 26,300
increases by 26,300
3) units = 12800*2 = 25600 units ; selling price = 30*90%=$27
Contribution Income statement
Sales (25600*27) 691200
Variable expense (25600*18) 460800
Contribution margin 230400
Fixed expenses (171,600+38000) 209,600
Net income 20,800
4) New contribution margin = 12-.70
11.3
BEP(units) = (total fixed cost+target profit)/contribution per unit
(171600+4400)/11.3
15575.22
Sales units 15,575
5)
CM ratio = contribution/sales
15/30
50.00%
BEP(units) = total fixed cost/contribution margin per unit
(171600+51000)/15
14840
BEP(dollars) = 222600/50%
445200
CM ratio 50%
Break even point in units 14840
Break even point in dollars 445200
20300
b) Not Automated Automated
total per unit % total per unit %
Sales 609000 30 100% 609000 30 100%
Variable expenses 365400 18 60% 304500 15 50%
Contribution margin 243600 12 40% 304500 15 50%
Fixed expenses 171,600 222,600
Net operating income 72,000 81,900
c) yes